# Generalized cryptographic semi-regular sequences: A variant of Fröberg conjecture and a simple complexity estimation for Gröbner basis computation

### v1

@@ 3.1 Proofs of Theorem 1 (1) and (2)
Suppose that $(f_{1},\ldots,f_{m})$ is generalized cryptographic semi-regular. Then, applying Lemma 2.2.8, we have the inequality (2.2.3), the right hand side of which is equal to $D_{\rm new}$ for $m\geq n$. It also follows from Lemma 3.1.2 that $\mathrm{max.GB.deg}_{\prec}(I)$ is upper-bounded by $\sum_{j=1}^{n-1}(d_{j}-1)+1$, which is equal to $D_{\rm new}$ for $m=n-1$, as desired.∎
3.2 Proof of Proposition 1
As in the previous subsection, let $\prec$ be the graded reverse lexicographic ordering on $R=K[x_{1},\ldots,x_{n}]$ with $x_{n}\prec\cdots\prec x_{1}$. We also denote the restriction of $\prec$ on $K[x_{1},\ldots,x_{n-1}]$ by the same notation $\prec$. In the following, we shall prove Proposition 1, namely the inequality (1.0.1) becomes an equality if $\langle\mathrm{LM}_{\prec}(I)\rangle_{R}$ is weakly reverse lexicographic. Here, a weakly reverse lexicographic ideal is a monomial ideal $J$ such that if $x^{\alpha}$ is one of the minimal generators of $J$ then every monomial of the same degree which preceeds $x^{\alpha}$ must belong to $J$ as well (see also e.g., [34, Section 4] for the definition).
$I^{\prime}:=\langle I,x_{n}\rangle_{R}\cap K[x_{1},\ldots,x_{n-1}]=\langle(f_{1})|_{x_{n}=0},\ldots,(f_{m})|_{x_{n}=0}\rangle_{K[x_{1},\ldots,x_{n-1}]}.$
Since $R/\langle I,x_{n}\rangle$ and $K[x_{1},\ldots,x_{n-1}]/I^{\prime}$ are isomorphic as graded rings, they have the same Hilbert series, and $d_{\rm reg}(I^{\prime})=d_{\rm reg}(\langle I,x_{n}\rangle_{R})$. Put $D:=d_{\rm reg}(I^{\prime})=d_{\rm reg}(\langle I,x_{n}\rangle_{R})$. By comparing Hilbert series, it can be also easily checked that, the sequence $(f_{1},\ldots,f_{m},x_{n})\in R^{m+1}$ is cryptographic semi-regular if and only if $(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})\in K[x_{1},\ldots,x_{n-1}]^{m}$ is cryptographic semi-regular, equivalently these sequences are both $D$-regular.
Let $G$ (resp. $G^{\prime}$) be the reduced Gröbner basis of $I$ (resp. $I^{\prime}$) with respect to $\prec$. Note that, since $\langle I,x_{n}\rangle_{R}=\langle I^{\prime},x_{n}\rangle_{R}$, the reduced Gröbner basis of $\langle I,x_{n}\rangle_{R}$ is given by $G^{\prime}\cup\{x_{n}\}$, whence $\mathrm{max.GB.deg}_{\prec}(I^{\prime})=\mathrm{max.GB.deg}_{\prec}(\langle I,x_{n}\rangle_{R})$.
To prove Proposition 1, we start with extending some results in [28] to our case.
Lemma 3.2.1 (cf. [28, Corollary 2]).
With notation as above, we assume that $(f_{1},\ldots,f_{m},x_{n})\in R^{m+1}$ is cryptographic semi-regular (i.e., $(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})\in K[x_{1},\ldots,x_{n-1}]^{m}$ is cryptographic semi-regular), namely both the sequences are $D$-regular with $D:=d_{\rm reg}(I^{\prime})=d_{\rm reg}(\langle I,x_{n}\rangle_{R})$. Then, we have
for any $d$ with $d<D$.
Proof.
Since $(f_{1},\ldots,f_{m},x_{n})$ is $D$-regular, it follows from [14, Proposition 2 a)] (cf. [28, Corollary 1]), its sub-sequence $(f_{1},\ldots,f_{m})$ is also $D$-regular. Hence, we have
$\equiv{\rm HS}_{R/\langle I,x_{n}\rangle}(z)\cdot\frac{1}{1-z}\equiv{\rm HS}_{R/\langle I,x_{n}\rangle}(z)\cdot(1+z+z^{2}+\cdots)$
Since $(JR)_{d}=x_{n}^{d}J_{0}\oplus x_{n}^{d-1}J_{1}\oplus\cdots\oplus x_{n}J_{d-1}\oplus J_{d}$ for any of $J=I^{\prime}$ and $J=K[x_{1},\ldots,x_{n-1}]$ (note that this holds also for $d\geq D$), the right hand side of (3.2.2) is equal to $\dim_{K}R_{d}-\dim_{K}(I^{\prime}R)_{d}$, so that $\dim_{K}(I^{\prime}R)_{d}=\dim_{K}I_{d}$. Here, it follows from [22, Theorem 5.2.6] that $\dim_{K}(\langle\mathrm{LM}_{\prec}(I)\rangle_{R})_{d}=\dim_{K}(\langle\mathrm{LM}_{\prec}(I^{\prime}R)\rangle_{R})_{d}$. Since $\langle\mathrm{LM}_{\prec}(I^{\prime}R)\rangle_{R}=\langle\mathrm{LM}_{\prec}(I^{\prime})\rangle_{R}$, we finally obtain (3.2.1). ∎
Lemma 3.2.2 (cf. [28, Lemma 2]).
Under the same setting as in Lemma 3.2.1, for each degree $d<D$, we have
${\rm LM}_{\prec}(G)_{d}={\rm LM}_{\prec}(G^{\prime})_{d}.$ (3.2.3)
Proof.
The equality (3.2.3) can be proved by the induction on $d$, as follows. The case $d=0$ is trivial. Suppose that (3.2.3) holds for any $d$ with $10\leq d<D-$ 1.
Let $t$ be any monomial in ${\rm LM}_{\prec}(G)_{d+1}$, so that $t=\mathrm{LM}_{\prec}(g)$ for some $g\in G$. By $d+1<D$, it follows from Lemma 3.2.1 that
$(\langle\mathrm{LM}_{\prec}(I)\rangle_{R})_{d+1}=(\langle\mathrm{LM}_{\prec}(I^{\prime})\rangle_{R})_{d+1},$
whence ${\rm LM}_{\prec}(g)$ is divided by ${\rm LM}_{\prec}(g^{\prime})$ for some $g^{\prime}\in G^{\prime}$. Since $G$ is reduced, the induction hypothesis implies that $\mathrm{deg}(g^{\prime})=d+1$, and thus $\mathrm{LM}_{\prec}(g)=\mathrm{LM}_{\prec}(g^{\prime})$. Therefore, we obtain ${\rm LM}_{\prec}(G)_{d+1}\subset{\rm LM}_{\prec}(G^{\prime})_{d+1}$. The same argument shows that ${\rm LM}_{\prec}(G)_{d+1}\supset{\rm LM}_{\prec}(G^{\prime})_{d+1}$, as desired. ∎
Lemma 3.2.3.
Under the same setting as in Lemma 3.2.1, we also suppose that $R/\langle I,x_{n}\rangle$ is Artinian (equivalently $K[x_{1},\ldots,x_{n-1}]/I^{\prime}$ is Artinian), namely $D<\infty$. Furthermore, we assume that $\langle\mathrm{LM}_{\prec}(I)\rangle_{R}$ is weakly reverse lexicographic. Then we have $\mathrm{max.GB.deg}_{\prec}(I)\geq D$ and $\mathrm{max.GB.deg}_{\prec}(I^{\prime})=D$.
Proof.
Put $d:=\mathrm{max.GB.deg}_{\prec}(I)$. Assume for a contradiction that $d<D$. Since $\mathrm{Krull.dim}(R/\langle I,x_{n}\rangle)=0$, it follows from Lemma 2.1.5 that there is an element $g$ in $G$ with ${\rm LM}_{\prec}(g)=x_{n-1}^{d^{\prime}}$ for some $d^{\prime}\leq d$. Then, any monomial $t\in R_{d^{\prime}}$ with $t\succeq x_{n-1}^{d^{\prime}}$ (namely $t$ is any monomial in $K[x_{1},\ldots,x_{n-1}]_{d^{\prime}}$) belongs to $\langle{\rm LM}_{\prec}(I)\rangle_{R}=\langle{\rm LM}_{\prec}(G)\rangle_{R}$ by the weakly reverse lexicographicness. Here, by $d<D$ together with (3.2.3), we have $\langle{\rm LM}_{\prec}(G)\rangle_{R}=\langle{\rm LM}_{\prec}(G^{\prime})_{\leq d}\rangle_{R}$. This implies that $t$ also belongs to $\langle{\rm LM}_{\prec}(G^{\prime})\rangle_{K[x_{1},\ldots,x_{n-1}]}$, and hence $K[x_{1},\ldots,x_{n-1}]_{d^{\prime}}/I^{\prime}_{d^{\prime}}=0$. Thus, we have $d\geq d^{\prime}\geq d_{\rm reg}(I^{\prime})$, a contradiction.
By the same argument as above, we can show that $\mathrm{max.GB.deg}_{\prec}(I^{\prime})=d_{\rm reg}(I^{\prime})$ as follows. Note that $\mathrm{max.GB.deg}_{\prec}(I^{\prime})\leq d_{\rm reg}(I^{\prime})$ follows from a general fact written in the second paragraph of Remark 3.2.4 below. Thus, we assume to the contrary that $d:=\mathrm{max.GB.deg}_{\prec}(I^{\prime})<D:=d_{\rm reg}(I^{\prime})$. Then, it follows from (3.2.3) that ${\rm LM}_{\prec}(G^{\prime})={\rm LM}_{\prec}(G)_{\leq d}$ and the ideal $\langle{\rm LM}_{\prec}(G^{\prime})\rangle_{K[x_{1},\ldots,x_{n-1}]}$ has the weak reverse lexicographicness up to $d$. Since there is an element $g$ in $G^{\prime}$ with ${\rm LM}_{\prec}(g)=x_{n-1}^{d^{\prime}}$ for some $d^{\prime}\leq d$, any monomial in $K[x_{1},\ldots,x_{n-1}]_{d^{\prime}}$ belongs to $\langle{\rm LM}_{\prec}(G^{\prime})\rangle_{K[x_{1},\ldots,x_{n-1}]}$ and so $K[x_{1},\ldots,x_{n-1}]_{d^{\prime}}/I^{\prime}_{d^{\prime}}=0$, which implies $d\geq d^{\prime}\geq D$, a contradiction. ∎
Remark 3.2.4.
In Lemma 3.2.3, when $R/\langle I,x_{n}\rangle$ is Artinian (or equivalently $K[x_{1},\ldots,x_{n-1}]/I^{\prime}$ is Artinian), we can easily prove $\mathrm{max.GB.deg}_{\prec}(I^{\prime})=D$ if $\langle\mathrm{LM}_{\prec}(I^{\prime})\rangle_{K[x_{1},\ldots,x_{n-1}]}$ is a weakly reverse lexicographic ideal, assuming neither that $(f_{1},\ldots,f_{m},x_{n})$ is cryptographic semi-regular nor that $\langle\mathrm{LM}_{\prec}(I)\rangle_{R}$ is weakly reverse lexicographic.
In general, for any homogeneous ideal $J\subset R$ with $\mathrm{Krull.dim}(R/J)=0$, we have $\mathrm{max.GB.deg}_{\prec}(J)\leq\mathrm{reg}(J)=d_{\rm reg}(J)$ by [4, Corollary 2.5] (see also [23, Proposition 4.15]), and it is straightforward that the equality holds if $\langle\mathrm{LM}_{\prec}(J)\rangle_{R}$ is a weakly reverse lexicographic.
Proof of Proposition 1:
Put $D:=d_{\rm reg}(\langle I,x_{n}\rangle_{R})=d_{\rm reg}(I^{\prime})$ and $D^{\prime}:=\widetilde{d}_{\rm reg}(I)$. From our assumption that $R/\langle I,x_{n}\rangle$ is Artinian, we have the inequality (1.0.1), i.e., $\mathrm{max.GB.deg}(I)\leq\mathrm{max}\{D,D^{\prime}\}$. The case where $\mathrm{Krull.dim}(R/I)=0$ is clear from Remark 3.2.4, so we assume that $\mathrm{Krull.dim}(R/I)=1$. Note that, for the reduced Gröbner basis $G$ of $I$, the elements of $\mathrm{LM}_{\prec}(G)$ are the minimal generators of the monomial ideal $\langle\mathrm{LM}_{\prec}(I)\rangle_{R}=\langle\mathrm{LM}_{\prec}(G)\rangle_{R}$.
First we consider the case where $D^{\prime}\geq D$, so that $\max\{D,D^{\prime}\}=D^{\prime}$. Put $d:=\mathrm{max.GB.deg}_{\prec}(I)$. Then, it follows from Macaulay’s basis theorem (cf. [26, Theorem 1.5.7]) that $R_{d}/I_{d}$ has a $K$-linear basis of the form
which is called the standard monomial basis. We set $r:=\dim_{K}R_{d}/I_{d}$, and write this standard monomial basis as $\{t_{1},\ldots,t_{r}\}$. In the following, we prove that $\{t_{1}x_{n}^{s},\ldots,t_{r}x_{n}^{s}\}$ is a $K$-linear basis of $R_{d+s}/I_{d+s}$ for any $s$ with $s\geq 1$, from which we have $\dim_{K}R_{d}/I_{d}=\dim_{K}R_{d+s}/I_{d+s}$ for any positive integer $s$, so that $d\geq D^{\prime}$ and therefore $d=D^{\prime}$.
By Lemma 3.2.3, we have $d\geq D$, and thus it can be shown similarly to Proof of Theorem 1 (1) that the multiplication-by-$x_{n}^{s}$ map from $R_{d}/I_{d}$ to $R_{d+s}/I_{d+s}$ is surjective (since $R/\langle I,x_{n}\rangle$ is Artinian). Therefore $B_{s}:=\{t_{1}x_{n}^{s},\ldots,t_{r}x_{n}^{s}\}$ generates $R_{d+s}/I_{d+s}$. Suppose to contrary that $B_{s}$ is not a basis. In this case, there exists an $i$ such that $t_{i}x_{n}^{s}$ is divisible by $\mathrm{LM}_{\prec}(g)$ for some $g\in G$. Indeed, since the elements of $B_{s}$ are linearly dependent over $K$, we obtain $f:=a_{1}t_{1}x_{n}^{s}+\cdots+a_{r}t_{r}x_{n}^{s}\in I_{d+s}$ for some $(a_{1},\ldots,a_{r})\in K^{r}\smallsetminus\{(0,\ldots,0)\}$. Thus, for the least $i$ with $a_{i}\neq 0$, the leading monomial of $f$ is $t_{i}x_{n}^{s}$, which is divisible by $\mathrm{LM}_{\prec}(g)$ for some $g\in G$. Putting $u_{i}=\mathrm{GCD}(t_{i},\mathrm{LM}_{\prec}(g))$ and $s_{i}=t_{i}/u_{i}$, we have $u_{i}s_{i}x_{n}^{s}=t_{i}x_{n}^{s}$. Since $t_{i}=s_{i}u_{i}$ is not divisible by $\mathrm{LM}_{\prec}(g)$, we can write $u_{i}x_{n}^{s^{\prime}}=\mathrm{LM}_{\prec}(g)$ for some $s^{\prime}\leq s$. (We note that $\deg s_{i}=d-\deg u_{i}\geq\deg{\rm LM}_{\prec}(g)-\deg u_{i}=s^{\prime}$.) Note that $s^{\prime}\geq 1$ since otherwise $t_{i}$ is divisible by $\mathrm{LM}_{\prec}(g)$.
Take an arbitrary monomial $s^{\prime}_{i}$ such that $\deg s^{\prime}_{i}=s^{\prime}$ and $s_{i}^{\prime}$ divides $s_{i}$. Then, by the weakly reverse lexicographicness, as ${\rm LM}_{\prec}(g)=u_{i}x_{n}^{s^{\prime}}\preceq u_{i}s^{\prime}_{i}$, the monomial $s^{\prime}_{i}u_{i}$ should belong to $\langle{\rm LM}_{\prec}(G)\rangle_{R}$. Moreover, since $s^{\prime}_{i}u_{i}$ divides $t_{i}=s_{i}u_{i}$, the monomial $t_{i}=s_{i}u_{i}$ also belongs to $\langle{\rm LM}_{\prec}(G)\rangle_{R}$, which is a contradiction.
3.3 Proof of Theorem 1 (3): Complexity estimation
@@
We shall prove (3) of Theorem 1. For this, we estimate the complexity of computing a Gröbner basis of a homogeneous ideal with Macaulay matrices.
@@
Let $F=\{f_{1},\ldots,f_{m}\}\subset R$ be a set of homogeneous polynomials in $R\smallsetminus K$, put $d_{j}:=\deg(f_{j})$ for $1\leq j\leq m$, and let $\prec$ be an arbitrary graded monomial ordering on $R$. The complexity of the Gröbner basis computation for $F$ is estimated as that of computing the reduced row echelon form (RREF) of the degree-$D$ Macaulay matrix, denoted by $M_{\leq D}(F)$, where $D$ is an arbitrary upper-bound on $\mathrm{max.GB.deg}_{\prec}(F)$. Here, the degree-$D$ Macaulay matrix $M_{\leq D}(F)$ (in homogeneous case) is a block matrix of the form
@@
with $d_{0}:=\min\{\deg(f_{j}):1\leq j\leq m\}$, where each $M_{d}(F)$ is a homogeneous Macaulay matrix defined as follows: For each $d$ with $d\geq d_{0}$, the degree-$d$ homogeneous Macaulay matrix $M_{d}(F)$ of $F$ has columns indexed by the terms of $R_{d}$ sorted, from left to right, according to the chosen order $\prec$. The rows of $M_{d}(F)$ are indexed by the polynomials $m_{i,j}f_{j}$, where $m_{i,j}\in R$ is a term such that $\deg(m_{i,j}f_{j})=d$.
@@
    [2, Proposition 1] The RREF of $M_{\leq D}(F)$ is computed in
@@
    Hence, if $\prec$ is a graded order, then the complexity of the Gröbner basis computation for $F$ is upper-bounded by (3.3.2) for $D=\mathrm{max.GB.deg}_{\prec}(F)$.
@@
    Hence, if $\prec$ is a graded order, and if $\mathrm{max.GB.deg}_{\prec}(F)=D=O(n)$, then the complexity of the Gröbner basis computation for $F$ is upper-bounded by (3.3.2).
@@
where we put $k:=D-d$. It follows from $D=O(n)$ that $c$ is upper-bounded by a constant. By $10<\varepsilon<$ 1, it is clear that $\frac{c+\varepsilon}{c+1}$ is also upper-bounded by a positive constant smaller than $1$, whence
$\frac{1}{1-\left(\frac{c+\varepsilon}{c+1}\right)^{\omega}}=O(1)$
as $n\to\infty$. We have proved the assertion (2). ∎

### v2

@@ 2.3 Genericness and Fröberg’s conjecture
whose right hand side is upper-bounded by $D^{(n,m)}$ for $m\geq n$. On the other hand, since $(f_{1},\ldots,f_{m})$ is generalized cryptographic semi-regular, it follows from Lemma 2.2.11 that we have the inequality (2.2.3), the right hand side of which is nothing but $D^{(n,m)}$ for $m\geq n$. Therefore, we obtain (1.3).
If $m=n-1$, it follows from Lemma 3.1.1 that $\mathrm{max}\{d_{\rm reg}(I+\langle x_{n}\rangle),\widetilde{d}_{\rm reg}(I)\}$ is upper-bounded by $\sum_{j=1}^{n-1}(d_{j}-1)+1$, which is equal to $D^{(n,m)}$ for $m=n-1$.∎
@@
Next, we shall prove (2) of Theorem 1.1. For this, we estimate the complexity of computing a Gröbner basis of a homogeneous ideal with Macaulay matrices.
@@
Let $\prec$ be an arbitrary graded monomial ordering on $R$. It is well-known that the complexity of the Gröbner basis computation for the homogeneous set $F$ is estimated as that of computing the reduced row echelon form (RREF) of the degree-$D$ Macaulay matrix, denoted by $M_{\leq D}(F)$, where $D$ is an arbitrary upper-bound on $\mathrm{max.GB.deg}_{\prec}(F)$. Here, the degree-$D$ Macaulay matrix $M_{\leq D}(F)$ (in homogeneous case) is a block matrix of the form
@@
with $d_{0}:=\min\{\deg(f_{j}):1\leq j\leq m\}$, where each $M_{d}(F)$ is a homogeneous Macaulay matrix defined as follows: For each $d$ with $d\geq d_{0}$, the degree-$d$ homogeneous Macaulay matrix $M_{d}(F)$ of $F$ has columns indexed by the monomials of $R_{d}$ sorted, from left to right, according to the chosen order $\prec$. The rows of $M_{d}(F)$ are indexed by the polynomials $m_{i,j}f_{j}$, where $m_{i,j}\in R$ is a monomial such that $\deg(m_{i,j}f_{j})=d$. For a monomial $t\in\mathcal{T}_{d}$ of degree $d$, the $(m_{i,j}f_{j},t)$-entry of $M_{d}(F)$ is defined as the coefficient of $t$ in $m_{i,j}f_{j}$.
@@
    [3, Proposition 1] The RREF of $M_{\leq D}(F)$ is computed in
@@
    Hence, if $\prec$ is a graded ordering, then the complexity of the Gröbner basis computation for $F$ is upper-bounded by (3.1.1) for $D=\mathrm{max.GB.deg}_{\prec}(F)$.
@@
    Hence, if $\prec$ is a graded ordering, and if $\mathrm{max.GB.deg}_{\prec}(F)=D=O(n)$, then the complexity of the Gröbner basis computation for $F$ is upper-bounded by (3.1.2).
@@
where we put $k:=D-d$. It follows from $D=O(n)$ that $c$ is upper-bounded by a constant. By $10<\varepsilon<$ 1, it is clear that $\frac{c+\varepsilon}{c+1}$ is also upper-bounded by a positive constant smaller than $1$, whence $1/{(1-(\frac{c+\varepsilon}{c+1})^{\omega})}=O(1)$ as $n\to\infty$. We have proved the assertion (2). ∎